Find the area enclosed by the curve r(θ)=1+2sinθ and the rays θ=0 and θ=π3
1
Expert's answer
2020-05-31T19:11:47-0400

area of the curve f is :

"\intop_0^{\frac{\pi}{3}} \frac{1}{2}{(f(\theta))}^2d\theta"


here, "f=1+2sin\theta"


"=\intop_0^{\frac{\pi}{3}} \frac{1}{2}{(1+2sin\theta)}^2d\theta"


"=\intop_0^{\frac{\pi}{3}} \frac{1}{2}{(1+4sin^2\theta+4sin\theta)}d\theta"


"=\frac{1}{2} [\frac{\pi}{3}-4cos\theta|_0^{\frac{\pi}{3}}+4 |\frac{\theta}{2}-\frac{sin2\theta}{4}|_0^{\frac{\pi}{3}}]"


because "\int sin^2 x dx=\frac{x}{2}-\frac{sin2x}{4}"


"= \frac{1}{2}[\frac{\pi}{3}+4.\frac{1}{2}+4(\frac{\pi}{6}-\frac{\sqrt3}{2})]"


"=\frac{\pi}{2}+1-\frac{\sqrt3}{4}"